OpenStudy (anonymous):

Using the differential equation dy/dx = 7 − y, find the intervals on the y-axis on which a nonconstant solution y = (x) is increasing. Find the intervals on the y-axis on which y = (x) is decreasing. PLEASE HELP!! Thanks!

OpenStudy (anonymous):

I m not sure if i am right.. but also quite sure :P first of all u do integration on both the left hand side and right hand side of the equation so u will get: \[\int\limits_{}^{}dy + y = \int\limits_{}^{} 7 dx\] \[y+ \frac{ y^2 }{ 2 } = 7x + c (constant)\]

OpenStudy (anonymous):

from there onwards u shud can figure out